# Is Long Context All You Need? Leveraging LLM’s Extended Context for NL2SQL Experiments, Analysis and Benchmark Paper

### v1

@@ 6.3. Long context in production and limitations
Our study reveals (Table 8) that a good chunk of requests (roughly 68%) from BIRD dev can reach  400k accumulated tokens and take up to  130 $T$ units. This would likely be prohibitive for most production services. And the cost will be even higher with techniques like self-consistency. Long context is a great means to add more appropriate information, but it has to be used with some care. To leverage long context properly in production, one may pick and choose which extra information and techniques are appropriate based on our analysis. For instance, one can skip disambiguation and online example generation to bound the latency of the majority of requests to at around ¡1.5 normalized seconds. The example can be generated offline and in Section 4.5 we show that example selection works well with synthetic examples. Using flash can further reduce the cost, roughly -20% in latency and -94% in dollars per request given the current pricing and the observed latency.
One limitation of our study, as we study the impact of long context and extra contextual information on NL2SQL, is missing benchmark datasets that model certain enterprise use cases where there are hundreds of very wide tables with thousands of columns. Also there is this aspect of less descriptive and similar table and column names that render schema linking much more challenging. Such use cases would be challenging in general for schema linking and also exacerbates the costs of long context techniques as with increased amounts of schema information. To this end, we tried to incorporate as many challenging NL2SQL benchmark datasets and evaluated how the long-context model perform with very low density information (e.g., providing full schema details and sample values from tens of irrelevant tables and DBs, while only a few columns are relevant, Table 3). We observe that extra information is less distracting and the long context techniques perform as good as the baselines without them. Filtering and ranking the most relevant information should still be critical in terms of both accuracy and the costs, but unless the retrieval is very accurate, providing more information for higher recall at the cost of lower precision is a good strategy. We expect these insights would hold for more extreme use cases, and become more evident as long context LLMs improves in the future; for instance, our long-context NL2SQL pipeline yields much stronger results on a enterprise data warehouse benchmark BEAVER than the baseline results from (Chen et al., 2024) (see Table 2).
@@ 7. Conclusion
In this work, we explored the potential of leveraging the extended context window offered by Google’s gemini-1.5-pro for NL2SQL. Our findings demonstrate that long-context LLMs can effectively utilize the additional context, achieving strong performances on various benchmarks (Table 2), including of 67.41% on the BIRD benchmark without fine-tuning or computationally expensive self-consistency techniques. This performance highlights the robustness of the long-context models in retrieving and reasoning over extensive contextual information. Specifically, we explore the performance implications of including entire DB schema details, user provided hints, sample column values, synthetic examples for many-shot ICL, for accurate SQL generation. We also show that self-correction and verification can improve accuracy further at the cost of increased latency and accumulated tokens. Overall, we observe that gemini-1.5-pro exhibits a strong retrieval capability over the extended context window, even with mostly irrelevant information. We also see that the long context model is more robust to or no longer “lost in the middle” (Liu et al., 2024a)).
Our study is a unique example of how enhanced capabilities of LLMs, in this case the extended large context size, impact the way we approach the NL2SQL problem. To the contrary to prior works where we focused on squeezing the filtered information onto limited context size, we explored the potential of providing more useful information to the model. And we showed that more information can be useful in NL2SQL, helping to overcome the semantic issues (SQL syntactic issues for the common SQL dialect are already well addressed by the base model). We also investigated the cost implication of using the long context techniques and conclude that it can be complementary and is more efficient with accurate schema and example retrievals and with a more emphasis on recall. In fact, perfect table schema retrieval would yield stronger performance (Appendix A.2) narrowing down the schema linking search space during generation (Chen et al., 2024; Floratou et al., 2024; Talaei et al., 2024), but it is very difficult to achieve in practice (Appendix A.1).
In a world where the retrieval and ranking are less than perfect, providing more information and relying on the long context models’ strong retrieval capability can be a viable alternative strategy, while being more expensive. Improving the cost efficiency of long context model serving would be an important area of research to make long context more practical.
@@ A.1. Table retrieval simulation
For each example in the benchmark, table retrieval operates as follows: 1) The simulated user view and query interactions narrow the search space down to  1-100 candidates; 2) The candidates are embedded and re-ranked with Gecko (Lee et al., 2024b) for relevance to the user query; 3) A fixed top-k candidates are passed on to the NL2SQL stage.
At the time of test, the recall of the first stage was 82%. As such, 82% is the maximum attainable end-to-end recall, and represents a near perfect result from the re-ranking stage.
@@ A.2. Impact of good column selection
We evaluate the impact of schema selection on execution accuracy using the BIRD dev. We compare a filtered schema against a perfect ground truth schema. The filtered schema, constructed similarly to CHESS (Talaei et al., 2024), incorporates relevant tables and columns, their descriptions, and relevant example values.  10 reports execution accuracy (Ex Acc), as well as table retrieval (TBR) and column retrieval (CR) recall and precision. While the ground truth schema achieves perfect performance (100%) on all metrics, the filtered schema demonstrates high recall, correctly identifying most relevant schema elements (97.69% table retrieval recall and 97.12% column retrieval recall). However, precision is lower (89.72% for table retrieval and 69.43% for column retrieval), indicating the inclusion of some irrelevant columns/tables. It is important to note that the schema filtering algorithm incurs substantial cost, requiring an average of 78 LLM calls and 339,965 tokens per request on the BIRD dev.
A.3. SQLite document retrieval
Here we show SQLite documentation retrieval results for example questions from BIRD dev dataset. We employed two chunking strategy to split the entire document by files and sections, respectively. We implemented a simple retrieval pipeline based on question & document similarity. The retrieval is challenging as the question contains limited information as to what documentation sections are relevant, and writing a SQL query can draw from many sections from the documentations. Even with the long context, we can only fit a select number of document chunks. We can summarize and try to compress to more chunks, but it will not be enough to fit the entire SQLite documentation or even all the relevant contents.
As shown in Table 6(b), having a large text from the documentation does not seem to affect the final generation quality much. Another interesting aspect of citing the SQLite documentation for NL2SQL is that the examples used in the documentation are often simple and generic as supposed to being specific to the user question and data. This is expected as the documentation tries to be illustrative of the general concepts and for the general audience.
Question: Among the cards with converted mana cost higher than 5 in the set Coldsnap, how many of them have unknown power?
SELECT SUM(CASE WHEN T1.power LIKE ’*’ OR T1.power IS NULL THEN 1 ELSE 0 END) FROM cards AS T1 INNER JOIN sets AS T2 ON T2.code = T1.setCode WHERE T2.name = ’Coldsnap’ AND T1.convertedManaCost $>$ 5
Retrieved coarse-grained chunks: Status Parameters for prepared statements, Built-In Scalar SQL Functions, List Of SQLite Functions
Retrieved fine-grained chunks: Rowid Tables, Definitions, Additional Notes, Searching, Status Parameters for prepared statements, Intellectual Property, Example Uses Of Row Values, List Of Core Functions, Release Notes, signed-number
Figure 7. Chunked SQLite documentation chunks retrieval results (relevant chunks, irrelevant chunks) with BIRD dev questions.
A.4. Prompt templates
The following are the generation, correction and verification prompt templates.
You are a SQLite SQL expert.
You need to generate SQLite SQL query given a question in natural language. The database (”{db_name}”) structure is defined by the following table schemas (comments after ’–’ provide additional column descriptions).
Given the ”Table creation statements” and the ”Question”, you need understand the database and columns.
Consider the natural language question to SQL query ”Examples”.
Also consider some useful ”Hints” if provided.
###Table creation statements###
{schema}
###Examples###
{examples}
###Documentation###
{documentation}
###Question###
{question}
(Hints: {hints})
Now generate SQLite SQL query to answer the given ”Question”.
Output the SQL query string ONLY.
Figure 8. Prompt template for NL2SQL generation.
You are a SQLite SQL expert.
Someone had a question and they tried to run a SQL query to fetch the data for it. However, the query execution failed for some error. Now you need to fix the query based on the previous execution error.
The database structure is defined by the following table schemas (comments after ’–’ provide additional column descriptions).
###Table creation statements###
{schema}
The original question is:
{question}
The SQL query executed was:
{sql}
The execution failed with error:
{error}
Based on the question, table schemas and the errored query, analyze what the query was trying to achieve and fix the error.
If the error cannot be fixed by fixing the query, for example, connection error or permission error, just output the original query. Otherwise, think step by step about generating correct SQLite SQL result!
Analyze the error and how to fix.
DONT FORGET Additional rules to generate correct SQLite SQL dialect:
{rules}
When you are OK with the fixed query, output the sqlite query string ONLY. It should be the query in plain text. Your answer must be a single SQL statement.
Figure 9. Prompt template for fix & rewrite.
You are a SQLite SQL expert.
Your job is to verify the correctness of a given SQL query both syntactically and semantically. That means you would have to ensure the SQL query is syntatically correct, and also it does return what the user is asking for in the natural language ”Question.”
Given the ”Table creation statements” and the ”Question and Hints”, you need understand the database and the relevant table columns. The database structure is defined by the following table schemas (comments after ’–’ provide additional column descriptions).
Use all these information to identify the correct literals, table column names and also the correct join path.
Also pay close attention to the ”Hints” as they contain very important information to answer the question correctly.
###SQL to verify###
{sql}
###Question and Hints###
{question}
###Table Creation Statements###
{schema}
Remember, following hints very closely is the key to the correct answer!!!
If you think the SQL query is incorrect, then return an empty string ””. If you are confident that the SQL query is correct, return it as-is.
Figure 10. Prompt template for output verification.

### v3

@@ 6.3. Long context in production and limitations
Our study reveals (Table 8) that a good chunk of requests (roughly 68%) from BIRD dev can reach  400k accumulated tokens and take up to  130 $T$ units. This would likely be prohibitive for most production services. And the cost will be even higher with techniques like self-consistency. Long context is a great means to add more appropriate information, but it has to be used with some care. To leverage long context properly in production, one may pick and choose which extra information and techniques are appropriate based on our analysis. For instance, one can skip disambiguation and online example generation to bound the latency of the majority of requests to at around ¡1.5 normalized seconds. The example can be generated offline, and as shown in Section 4.5, example selection works well with synthetic examples. Using flash can further reduce the cost, roughly -20% in latency and -94% in dollars per request given the current pricing and the observed latency.
One limitation of our study, as we analyze the impact of long context and additional contextual information on NL2SQL, is the absence of benchmark datasets that accurately model enterprise use cases. Such cases typically involve hundreds of very wide tables with thousands of columns, as well as schemas with less descriptive and similar table and column names, making schema linking significantly more challenging. These scenarios not only pose inherent difficulties for schema linking, but further exacerbate the cost of long-context techniques due to the increased amount of schema information. To address this, we incorporated multiple challenging NL2SQL benchmark datasets and evaluated how the long-context model performs in cases with very low density information (e.g., providing full schema details and sample values from tens of irrelevant tables and DBs, while only a few columns are relevant, as shown in Table 3(b)). We observe that extra information is generally not distracting, and the long context techniques perform as well as the baselines without them. While filtering and ranking the most relevant information remain crucial for both accuracy and cost, our findings suggest that, unless retrieval is highly accurate, increasing recall by including more information - even at the expensive of lower precision - is a beneficial strategy. We expect that these insights will hold for more extreme use cases and become increasingly evident as long context LLMs continue to improve. For instance, our long-context NL2SQL pipeline yields significantly better results on the enterprise data warehouse benchmark BEAVER compared to the baseline results from (Chen et al., 2024) (see Table 2).
While this work focuses on evaluating the impact of providing various forms of contextual information directly within the LLM’s extended context window, we acknowledge that the explored information is not exhaustive. Notably, we did not investigate the utilization of knowledge graphs for enhancing NL2SQL performance. While it is expensive to build a comprehensive knowledge graph for a given domain, recent works (Allemang and Sequeda, 2024; Sequeda et al., 2024) have shown promising results using knowledge graph-based question answering over enterprise datasets. In future work, we are interested in exploring the integration of additional contextual sources, such as business ontologies and knowledge graph, to further leverage the capabilities of long-context LLMs.
@@ 7. Conclusion
In this work, we explored the potential of leveraging the extended context window offered by Google’s gemini-1.5-pro for NL2SQL. Our findings demonstrate that long-context LLMs can effectively utilize the additional context, achieving strong performance across multiple benchmarks (Table 2), including 67.41% execution accuracy on the BIRD benchmark - without fine-tuning or computationally expensive self-consistency techniques. This performance highlights the robustness of the long-context models in retrieving and reasoning over extensive contextual information. Specifically, we analyze the impact of including entire DB schema details, user provided hints, sample column values, and synthetic examples for many-shot ICL to improve SQL generation accuracy. We also show that self-correction and verification can further enhance accuracy, although at the cost of increased latency and accumulated tokens. Overall, we observe that gemini-1.5-pro exhibits strong retrieval capabilities over the extended context window, even in the presence of irrelevant information. Additionally, our findings indicate that the long context models are more robust and do not suffer from the “lost in the middle” (Liu et al., 2024a)) issue.
Our study is a unique example of how enhanced capabilities of LLMs, in this case the extended large context size, impact the way we approach the NL2SQL problem. In contrary to prior works that focused on squeezing the filtered information into a limited context window, we explored the potential of providing additional useful information to the model. We showed that the additional information can be useful in NL2SQL by helping to resolve semantic issues (SQL syntactic issues for the common SQL dialect are already well addressed by the base model). Furthermore, we investigated the cost implications of using the long context techniques and concluded that they can be complementary and more efficient when combined with accurate schema and example retrievals, with an emphasis on recall. In fact, perfect table schema retrieval would yield stronger performance (Appendix A.2) by narrowing the schema linking search space during SQL generation (Chen et al., 2024; Floratou et al., 2024; Talaei et al., 2024). However, achieving perfect retrieval is highly challenging in practice (Appendix A.1).
In real-world scenarios where the retrieval and ranking are sub-optimal, providing more information and relying on the long context models’ strong retrieval capability can offer a viable, though more expensive, alternative strategy. Improving the cost efficiency of long context model serving would be an important area of research to make long context more practical.
@@ A.1. Table retrieval simulation
For each example in the benchmark, table retrieval operates as follows: 1) The simulated user view and query interactions narrow the search space down to  1-100 candidates; 2) The candidates are embedded and re-ranked with Gecko (Lee et al., 2024b) for relevance to the user query; 3) A fixed top-k candidates are passed on to the NL2SQL stage. At the time of test, the recall of the first stage was 82%. As such, 82% is the maximum attainable end-to-end recall, and represents a near perfect result from the re-ranking stage.
@@ A.2. Impact of good column selection
We evaluate the impact of schema selection on execution accuracy using the BIRD dev. We compare a filtered schema against a perfect ground truth schema. The filtered schema, constructed similarly to CHESS (Talaei et al., 2024), incorporates relevant tables and columns, their descriptions, and relevant example values.  12 reports execution accuracy (Ex Acc), as well as table retrieval (TBR) and column retrieval (CR) recall and precision. While the ground truth schema achieves perfect performance (100%) on all metrics, the filtered schema demonstrates high recall, correctly identifying most relevant schema elements (97.69% table retrieval recall and 97.12% column retrieval recall). However, precision is lower (89.72% for table retrieval and 69.43% for column retrieval), indicating the inclusion of some irrelevant columns/tables. It is important to note that the schema filtering algorithm incurs substantial cost, requiring an average of 78 LLM calls and 339,965 tokens per request on the BIRD dev.